OpenStudy (anonymous):

(sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2x

OpenStudy (psymon):

Proving the identity, right?

OpenStudy (anonymous):

Yeah

OpenStudy (psymon):

I would start with turning tangent and cotangent into sines and cosines.

OpenStudy (anonymous):

the sin and cos are separated

OpenStudy (psymon):

Yeah, I know what you mean with how you typed it. Alright, so cosines cancel on the left side of your parenthesis and the right becomes cos^2(x)/sinx. Think you notice what to do after that, or not sure?

OpenStudy (anonymous):

Do you think you could show me by like typing out because I think I know what to do but I just want to be certain ?

OpenStudy (psymon):

Sure

OpenStudy (psymon):

Sure\[sinx(sinx - \frac{ \cos ^{2}x }{ sinx }) = 1-2\cos ^{2}x\] So thats what it should look like.

OpenStudy (anonymous):

Ok then would you distribute the sin to the other sin on the left side

OpenStudy (psymon):

Yes, you would distribute sinx to both terms inside of the parenthesis.

OpenStudy (anonymous):

ok so it then would be sin^2 x- cosx /sin^2x

OpenStudy (psymon):

Sorry. And when you distribute sinx, itll cancel out in the second term:

OpenStudy (anonymous):

could you show me

OpenStudy (psymon):

\[sinx(sinx - \frac{ \cos ^{2x} }{ sinx }) = 1-2\cos ^{2}x\] This becomes: \[(\sin ^{2}x - \frac{ sinxcos ^{2}x }{ sinx }) = 1-2\cos ^{2}x\] Can you see from here?

OpenStudy (anonymous):

I think so

OpenStudy (psymon):

Awesome. So once sinx cancels on that right term, know that sin^2(x) + cos^2(x) = 1. Using that fact, you can do a substitution that will lead to the finality of your proof : ) I have to head out now unfortunately. Good luck ^_^